# Chlorophyll-a Mapping and Prediction in the Mar Menor Lagoon Using C2RCC-Processed Sentinel 2 Imagery

### v1

@@ V-A Results
Two dates were selected for illustration: 2022-07-14 and 2025-07-28. The first corresponds to the well-documented whiting event, when a white spot appeared on the western side of the lagoon near the Albujón seasonal watercourse. Such phenomena are usually linked to phytoplankton blooms or riverine sediments, although in this case the cause was uncertain. Previous reports [62, 63] noted consistently higher Chl-a concentrations within the white spot compared to surrounding areas. Accordingly, the predicted map for this date is expected to highlight the white spot with higher Chl-a values than its vicinity. The other date chosen is 2025-07-28, because recent monitoring reports indicated a potential eutrophication episode triggered by increasing Chl-a levels in the lagoon [64, 14]. Consequently, the output for this date is expected to show higher Chl-a concentrations than in July 2022, when average values were relatively low. Additionally, a general pattern observed is that Chl-a concentration tends to increase with depth [15, 62].
@@
Figure 3 shows the white spot on the western side, with Chl-a concentration increasing with depth. The concentrations at shallower depth range from 0.3 to 1.4 mg/m3, while at greater depths concentrations increase nearly to 3 mg/m3 in the white spot area. There are some parts in which the borders show much higher concentrations, but those values should be ignored since they are produced by soil reflection due to its shallow depth.
@@
These representations, together with additional examples provided in the Supplementary Material, confirm that the algorithms behave as expected. Further testing revealed that the KNN model applied at 2–3 m depth tends to overestimate Chl-a under extreme conditions, such as during the 2016 eutrophication crisis. While the other depths reported high Chl-a concentrations ranging from 10 to 20 mg/m3, this layer saturated the color scale, indicating that its predictions could be higher than expected. This discrepancy may be related to the fact that this layer relied on TOA reflectances without atmospheric correction. In contrast, the atmospheric correction applied with C2X-Complex at the other depths introduces a characteristic noise pattern, which can be observed in Figures 3 and 4.
VI Conclusions and future work
@@
This study presented an end-to-end methodology to predict and map Chl-a concentrations in the Mar Menor lagoon by integrating nearly a decade of Sentinel 2 imagery with in situ buoy measurements. The approach combined atmospheric correction using the C2RCC family of neural network processors, diverse multispectral band combinations, and multiple machine and DL models, evaluated within a robust cross-validation framework.
@@
The results demonstrate that the predictive performance varies with depth, processing method, and aggregation window. At the surface (0-1 m), the best performance was obtained with ‘C2X-Complex_rhow_9x9”, achieving an R2 of 0.89 with both XGBoost and the ensemble. For the 1-2 m layer, the strongest results were provided by “C2X-Complex_rhow_5x5”, with the ensemble reaching an R2 of 0.87 and RMSE of 1.53 mg/m³, closely followed by CatBoost. At 2-3 m, unprocessed TOA reflectances with 15x15 pixels aggregation outperformed atmospherically corrected datasets, with KNN delivering competitive results (R2 of 0.81, RMSE of 1.61 mg/m³). At the deepest layer (3-4 m), performance declined by nearly 0.2 in R2 compared to shallower depths, but “C2X-Complex_rhow_5x5” and Random Forest provided the most consistent results (R2 = 0.66, RMSE = 1.48 mg/m³). These findings underline the importance of carefully selecting processing variants and spatial aggregation strategies according to depth.
@@
Compared to previous approaches reported in the literature, the methodology developed in this study offers several distinctive advantages. While most existing works in the Mar Menor and other similar systems have relied on short time series, narrower ranges of chlorophyll-a values, or surface-only estimates, our framework integrates nearly a decade of Sentinel 2 imagery with in situ buoy measurements and explicitly extends predictions across the water column. By combining advanced atmospheric correction through C2RCC and its variants with a systematic evaluation of band combinations, aggregation strategies, and a wide array of ML and DL models, the proposed solution achieves robust results at multiple depths, which were validated by reproducing known events such as the 2016 eutrophication crisis and the 2022 whiting phenomenon. This ability to generate spatially explicit chlorophyll-a maps not only at the surface but also at subsurface layers provides a more comprehensive perspective of lagoon dynamics, delivering a richer and more informative output than existing empirical or single-depth approaches. Thus, this capability enhances monitoring by combining the temporal continuity of buoy measurements with the spatial coverage of remote sensing.
@@
In summary, the framework provides a reproducible and transferable approach for Chl-a estimation in optically complex waters. Its integration of Sentinel 2 imagery, advanced atmospheric correction, and a wide array of learning algorithms makes it a valuable tool for long-term monitoring of the Mar Menor and potentially other vulnerable coastal and inland water bodies.
@@
However, several limitations must be acknowledged. Model performance decreases with depth, reflecting the inherent challenges of retrieving reliable reflectances from optically complex waters, where signal attenuation and variability increase with distance from the surface. Cloud cover imposes an additional permanent constraint, reducing the temporal resolution of usable Sentinel 2 imagery. Moreover, the current processing workflow involves multiple sequential steps (downloading SAFE files, atmospheric correction, feature generation, aggregation, and model inference), which implies substantial effort for each new set of maps generated.
@@
  • Automating chlorophyll map generation by developing a pipeline that, starting from a selected date, processes Sentinel 2 SAFE files in SNAP, extracts reflectances for all pixels in the Mar Menor, incorporates additional features, and performs model inference to generate chlorophyll maps in a fully automated manner, addressing the third limitation previously mentioned.
  • Study Chl-a concentrations over the last decades by generating maps using the full historical Sentinel 2 archive and include them into the digital twin developed by [65].

### v2

@@ 5.1 Results
Two dates were selected for illustration: 2022-07-14 and 2025-07-28. The first corresponds to the well-documented whiting event, when a white spot appeared on the western side of the lagoon near the Albujón seasonal watercourse. Such phenomena are usually linked to phytoplankton blooms or riverine sediments, although in this case the cause was uncertain. Previous reports [Belich, CSICManchaBlanca] noted consistently higher Chl-a concentrations within the white spot compared to surrounding areas. Accordingly, the predicted map for this date is expected to highlight the white spot with higher Chl-a values than its vicinity. The other date chosen is 2025-07-28, because recent monitoring reports indicated a potential eutrophication episode triggered by increasing Chl-a levels in the lagoon [MarMenorAlerta, pressMarMenorEntra2025]. Consequently, the output for this date is expected to show higher Chl-a concentrations than in July 2022, when average values were relatively low. Additionally, a general pattern observed is that Chl-a concentration tends to increase with depth [erena2017analisis, Belich].
@@
Figure 3 shows the white spot on the western side, with Chl-a concentration increasing with depth. The concentrations at shallower depth range from 0.3 to 1.2 mg/m3, while at greater depths concentrations increase nearly to 3 mg/m3 in the white spot area.
@@
While in some cases unexpected high or low concentrations are observed near the shoreline, these values should be interpreted with caution. The integrated bathymetric lines confirm that these overestimations or underestimations, depends on the particular case, occur in extremely shallow areas, where bottom reflectance interference, uncovered by the low depth contours, likely biases the spectral signal.
The results and representations obtained, together with additional examples provided in the Supplementary Material, confirm that the algorithms behave as expected, and based on those, several technical recommendations are proposed for monitoring Chl-a in optically complex coastal systems. A depth-stratified processing strategy is recommended, where specialized neural-network processors are prioritized for surface layers while unprocessed TOA reflectances are considered for deeper segments to avoid potential spectral over-correction in low-signal environments. To mitigate sensor noise and single-pixel variability in high-resolution data, spatial aggregation using windows of at least $\times9$ pixels is essential for securing stable and representative estimates. Furthermore, the integration of seasonal categorical metadata with gradient-boosted ensembles, specifically CatBoost, is advised to effectively manage the high multicollinearity inherent in spectral bands and derived indices. Finally, a multi-objective optimization framework incorporating several metrics, sucha as RMSLE and R2, could be crucial for ensuring model robustness and sensitivity across a lagoon’s full trophic range.
5.2 Workflow automation
To ensure full reproducibility and streamline the chlorophyll mapping workflow, a Docker-based pipeline was developed to automate the complete process, from image retrieval to map generation. The containerized environment, built from an Ubuntu 22.04 base image, integrates all dependencies including Python 3.11.0, SNAP 12.0.0, and the required Python libraries. The system requires only a single input parameter—the target date in %Y-%m-%d format—to execute all processing stages automatically. Within the container, the main script run_pipeline.py orchestrates the workflow, sequentially invoking modules for Sentinel 2 product download, atmospheric correction, reflectance processing, model inference, and visualization. Intermediate products are stored in organized subdirectories, and the final Chl-a maps are saved in a persistent folder. Additionally, a complementary script, check_dates.py, allows users to evaluate cloud coverage in the area of interest for a specific date range before running the pipeline. The container can be executed with a single command, producing all outputs in approximately 10–15 minutes per date. This implementation enhances reproducibility, portability, and scalability, enabling consistent deployment across different computing environments.
6 Conclusions and future work
@@
This study presented an end-to-end methodology to predict and map Chl-a concentrations in the Mar Menor lagoon by integrating nearly a decade of Sentinel 2 imagery with in situ buoy measurements. The approach combined atmospheric correction using the C2RCC family of neural network processors, diverse multispectral band combinations, and multiple ML models, evaluated within a robust cross-validation framework. Then, a fully automated pipeline is provided for reproducibility and applicability purposes.
@@
The results demonstrate that the predictive performance varies with depth, processing method, and aggregation window. At the surface (0-1 m), the best performance was obtained with the ’C2X-Complex_rhow_15x15’ configuration using CatBoost, achieving an RMSLE below 0.34 and an $R^{2}$ above 0.75. For the 1-2 m layer, the strongest results were similarly provided by ’C2X-Complex_rhow_15x15’ with CatBoost, reaching an $R^{2}$ above 0.75 and a satisfactory RMSLE between 0.35 and 0.4. At 2-3 m, unprocessed TOA reflectances with $\times15$ pixel aggregation outperformed atmospherically corrected datasets, with CatBoost delivering an RMSLE of 0.38 and an $R^{2}$ near 0.70. At the deepest layer (3-4 m), performance declined compared to shallower depths, but ’TOA_15x15’ and CatBoost provided the most consistent results, yielding an $R^{2}$ near 0.60 and an RMSLE around 0.39. These findings underline the importance of carefully selecting processing variants and spatial aggregation strategies according to depth.
@@
Compared to previous approaches reported in the literature, the methodology developed in this study offers several distinctive advantages. While most existing works in the Mar Menor and other similar systems have relied on short time series, narrower ranges of Chl-a values, or surface-only estimates, our framework integrates nearly a decade of Sentinel 2 imagery with in situ buoy measurements and explicitly extends predictions across the water column. By combining advanced atmospheric correction through C2RCC and its variants with a systematic evaluation of band combinations, aggregation strategies, and a wide array of ML models, the proposed solution achieves robust results at multiple depths, which were validated by reproducing known events such as the 2016 eutrophication crisis and the 2022 whiting phenomenon. This ability to generate spatially explicit Chl-a maps not only at the surface but also at subsurface layers provides a more comprehensive perspective of lagoon dynamics, delivering a richer and more informative output than existing empirical or single-depth approaches. Thus, this capability enhances monitoring by combining the temporal continuity of buoy measurements with the spatial coverage of remote sensing.
@@
In summary, the framework provides a reproducible and transferable approach for Chl-a estimation in optically complex waters. Its integration of Sentinel 2 imagery, advanced atmospheric correction, and a wide array of learning algorithms makes it a valuable tool for long-term monitoring of the Mar Menor and potentially other vulnerable coastal and inland water bodies. Furthermore, the automated pipeline puts the framework available to any interested party in creating chlorophyll maps of the Mar Menor immediately.
@@
However, several limitations must be acknowledged. Model performance decreases with depth, reflecting the inherent challenges of retrieving reliable reflectances from optically complex waters, where signal attenuation and variability increase with distance from the surface. Moreover, cloud cover imposes an additional permanent constraint, reducing the temporal resolution of usable Sentinel 2 imagery.
Beyond environmental factors, the temporal synchronization between satellite overpasses (10:30-11:00 UTC) and in situ measurements (10:00-13:00) introduces a maximum offset of 2.5 hours. While sub-daily fluctuations remain a boundary condition, this temporal gap is within acceptable limits for robust regional monitoring.
Regarding the modeling paradigm, the “black-box” nature of the algorithms used limits the direct interpretability of the underlying ecological processes. While these models excel at capturing complex non-linear relationships for prediction, they do not explicitly model the specific bio-optical equations that govern light-water interactions.
Finally, model transferability remains a constraint. The weights and hyperparameters optimized in this study are specifically tailored to the unique hypersaline and turbid conditions of the Mar Menor. While the end-to-end methodology is transferable, the models themselves would require another training using local ground truth before being applied to other coastal or inland water bodies with different inherent optical properties.
@@
  • Study Chl-a concentrations over the last decades by generating maps using the full historical Sentinel 2 archive and include them into the digital twin developed by [yeAdvancingMarineDigital2024] and in other services as WMS.
@@
  • Explore the integration of Explainable AI (XAI) techniques, such as SHAP or LIME, to unravel the contribution of specific spectral bands to the model’s decisions, bridging the gap between ML and ecological understanding.
@@
7 Declaration of generative AI and AI-assisted technologies in the manuscript preparation process
During the preparation of this work the authors used Google Gemini and DeepL Write to check the grammar of the manuscript and improve the clarity of the more technical sections. After using these tools, the authors reviewed and edited the content as needed and take full responsibility for the content of the published article.